

********************************************************************
UARS WEEKLY STATUS REPORT - 06 Feb 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 58% of
each day's data amounting to nearly 480 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm)
resolution, Mg II, and high resolution monitoring scans
during this past week for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003149 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1993099. MEPS data files between
1993100 and 1993139 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 January 2004 (Orbit 67676, GMT Day 024) through
30 January 2004 (Orbit 67780, GMT Day 030).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 024/2200 67689 REBOOT
HALOE 025/1319 67699 POWER ON
PEM 026/1316 67714 ZEPS & NEPS TO DAY ONLY
The beta angle ranged between the angles of 11.6 degrees to 33.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 41 degrees on 05 February 2004 (DOY 036).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.29 to 1.59 degrees C. Additional
battery performance data for each battery on 23 January and 30 January
2004 are provided in Appendix D.
The clock error ranged from 10.4 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
SUSIM = 01
SOLSTICE = 01
HALOE = 02
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to prepare for the uplink of the 3-TDRS
patch on 04 February 2004.
MEETINGS
The FOT attended a UARS FSW code review on 27 January 2004 to discuss
the test results obtained by FSW. All test data showed that the patch
performs as expected and the uplink remains on schedule.
The FOT also attended the UARS 3-TDRS ORR on 29 January 2004 to
present the work done in preparation for the new 3-TDRS capability to
a review panel. The government did not find any problems that
warranted a delay of the uplink schedule and gave the FOT the go ahead
to proceed.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 February 2004 (DOY 052).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
The 3-TDRS uplink day is scheduled for 04 February 2004 (DOY 035).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 January 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67684 024/1322 04-030 SUSIM POWER ON; PART B
67689 024/2200 04-031 SOLSTICE REBOOT
67699 025/1319 04-032 HALOE POWER ON; PART 1
67702 025/1754 04-033 HALOE POWER ON; PART 2
67709 026/0553 04-034 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67714 026/1316 04-035 PEM ZEPS & NEPS DAY ONLY
67727 027/0953 04-036 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67763 029/1923 04-037 OBC SOFT DUMP OF OBC MEMORY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 January 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 January 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67709 026/0553 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67727 027/0953 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 January 2004
23 January 2004 - GMT Day 023 Beta = 07.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-44.8 +4.50/+3.87 +28.4/-12.8 26.1 1.03 19.6
3 +11.2/-33.6 +3.23/+2.60 +27.2/-12.8 26.1 1.02 19.0
30 January 2004 - GMT Day 030 Beta = 33.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-179.2 +5.79/+4.50 +28.4/-14.4 25.1 1.04 20.9
3 +11.2/-123.2 +5.14/+3.55 +28.0/-15.2 25.1 1.03 20.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 January 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-006 FOT 67684 024/1322 34874 HPA Failure
04-007 FOT 67688 024/2017 N/A Could Not Load
04-008 SOLSTICE 67688 024/2017 N/A SOLSTICE Detectors RED
04-001 FOT 67694 024/0556 30832 Generic Late Acquisition
04-001 FOT 67701 025/1610 30832 Generic Late Acquisition
04-001 FOT 67737 028/0219 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-006 FOT 67632 021/0233 N/A Eclipse Hung
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-001 FOT See List Jan 04 30832 Generic Late Acquisition
04-004 FOT 67589 018/0624 N/A IP NOCC switch problems
04-006 FOT 67684 024/1322 34874 HPA Failure
04-007 FOT 67688 024/2017 N/A Could Not Load
04-008 SOLSTICE 67688 024/2017 N/A SOLSTICE Detectors RED